OpenStudy (samsan9):

how to solve the tan 165 using the sum and difference formulas?

OpenStudy (anonymous):

tan(165) = tan(120+45) *degrees

OpenStudy (samsan9):

ok i got \[-\sqrt{3}+1\]

OpenStudy (samsan9):

\[1+\sqrt{3}(1)\]

OpenStudy (anonymous):

you should have gotten sqrt(3) -2

OpenStudy (samsan9):

is rhat the answer?

OpenStudy (anonymous):

yes

OpenStudy (samsan9):

so wait do we just use the box method?